I need help understanding this problem.
x+3y=15
-8x+3y=-12
So far I have done it 3 times ive got the (3,_) but cant understand how to get 4 using process of elimination

Answer by stanbon(75887)

x+3y=15
-8x+3y=-12
----
Subtract and solve for "x":
9x = 27
x = 3
---
Solve for "y":
x + 3y = 15
3 + 3y = 15
3y = 12
y = 4
=============
Solution: (3,4)
==================
Cheers,
Stan H.
=============